OpenStudy (anonymous):

(x+2)/(x+4) = 3/x When I solved for x, I got x=3 and 4. Am I correct?

jhonyy9 (jhonyy9):

ok but when there is x on denominator what is the first step ,what is the restriction what we need taking for x ?

jhonyy9 (jhonyy9):

do you know it ?

jhonyy9 (jhonyy9):

so when is a fraction undefined ?

OpenStudy (anonymous):

when the denominator = 0?

jhonyy9 (jhonyy9):

yes so than what will be the restriction value for x ?

jhonyy9 (jhonyy9):

can you calculi it ?

OpenStudy (anonymous):

I don't really understand...

jhonyy9 (jhonyy9):

you have said that a fraction is undefined when the denominator is equal zero so there are 3 fractions with 3 denominators with x so for these denominators dont being equal zero what values not can getting x how can you calculi these values ?

jhonyy9 (jhonyy9):

sorry not 3 just two

jhonyy9 (jhonyy9):

ok what are the denominators of these two fractions ?

OpenStudy (anonymous):

x+4 and x

jhonyy9 (jhonyy9):

yes so x+4 and x when will be equal zero ,when x is equal what value ?

OpenStudy (anonymous):

-4 and 0?

jhonyy9 (jhonyy9):

yes right so these are the restrictions for x ok ?

OpenStudy (anonymous):

Yeah, okay.

jhonyy9 (jhonyy9):

so than now (x+2)/(x+4) = 3/x because there are two fractions with two different denominators what you need to do ?

OpenStudy (anonymous):

find common denominators?

jhonyy9 (jhonyy9):

yes is right but you can solve it by different ,easy style so like ,,cross multiplie" yes ?

jhonyy9 (jhonyy9):

what you will get than ?

OpenStudy (anonymous):

Yeah. x^2 + 2x over 3x-12

jhonyy9 (jhonyy9):

(x+2)/(x+4) = 3/x cross multiplie how you get ,,over" ?

jhonyy9 (jhonyy9):

x+2 3 ---- = --- x+4 x cross multiplie what mean ?

OpenStudy (anonymous):

over is the fraction. (x^2+2x)/(3x-12)

jhonyy9 (jhonyy9):

yes i understand it right but not is correct so what mean cross multiplie when there is x+2 3 ---- = --- x+4 x ...???

OpenStudy (anonymous):

You multiply x+2 and x. Then you multiply x+4 and 3. Right?

jhonyy9 (jhonyy9):

yes ,please write these here now

jhonyy9 (jhonyy9):

x(x+2)=3(x+4) yes ? than make the calculi

OpenStudy (anonymous):

then it would equal x^2+2x = 3x - 12

jhonyy9 (jhonyy9):

not is right

OpenStudy (anonymous):

What did I do wrong?

jhonyy9 (jhonyy9):

check it step by step term by term

jhonyy9 (jhonyy9):

x(x+2)=3(x+4) x(x+2) = ?

jhonyy9 (jhonyy9):

comen i like make you giving where not is right

jhonyy9 (jhonyy9):

how you have got on the right side - 12 ?

OpenStudy (anonymous):

Distributive property. 3(x-4). 3(x) -3(4).